# Demystifying Catastrophic Forgetting in Two-Stage Incremental Object Detector

### v1

@@ Abstract
Catastrophic forgetting is a critical chanllenge for incremental object detection (IOD). Most existing methods treat the detector monolithically, relying on instance replay or knowledge distillation without analyzing component-specific forgetting. Through dissection of Faster R-CNN, we reveal a key insight: Catastrophic forgetting is predominantly localized to the RoI Head classifier, while regressors retain robustness across incremental stages. This finding challenges conventional assumptions, motivating us to develop a framework termed NSGP-RePRE. Regional Prototype Replay (RePRE) mitigates classifier forgetting via replay of two types of prototypes: coarse prototypes represent class-wise semantic centers of RoI features, while fine-grained prototypes model intra-class variations. Null Space Gradient Projection (NSGP) is further introduced to eliminate prototype-feature misalignment by updating the feature extractor in directions orthogonal to subspace of old inputs via gradient projection, aligning RePRE with incremental learning dynamics. Our simple yet effective design allows NSGP-RePRE to achieve state-of-the-art performance on the Pascal VOC and MS COCO datasets under various settings. Our work not only advances IOD methodology but also provide pivotal insights for catastrophic forgetting mitigation in IOD. Code will be available soon.
@@ 2.1 Incremental Learning for Classification
Most influential incremental learning studies have focused on classification tasks. These methods can be categorized into regularization-based, structure-based, and replay-based methods. Some regularization-based methods enforce the stability of logits (Li & Hoiem, 2017; Zhang et al., 2023) or intermediate features (Simon et al., 2021) to preserve the learned knowledge, while others apply restrictions on the weight of the model (Kirkpatrick et al., 2017) or on gradients during optimization (Lopez-Paz & Ranzato, 2017; Wang et al., 2021). Structure-based methods are dedicated to learning specific parameters for different tasks, with a dynamically expanding architecture (Rusu et al., 2016) or grouped parameters in a static model (Fernando et al., 2017). For replay-based methods, they can be divided into experience replay (Buzzega et al., 2020; Zhu et al., 2021; Kong et al., 2023) and generative replay methods (Zhai et al., 2020; Kemker & Kanan, 2017), depending on the examples stored in a buffer or generated with a model. Recently, incremental learning based on foundation models such as CLIP (Radford et al., 2021) has also attracted attention. Research works such as L2P (Wang et al., 2022), O-LoRA (Wang et al., 2023a), and VPT-NSP2 (Lu et al., 2024) attempt to learn continuously based on the parameter-efficient transfer learning technique (Zhou et al., 2022; Jia et al., 2022; Xing et al., 2023) have achieved superior performance.
@@ 6 Conclusion
Similar to previous works, we use the Faster R-CNN architecture with a Resnet-50 (He et al., 2016) backbone pre-trained in ImageNet (Russakovsky et al., 2015). On PASCAL VOC dataset, we train the network with SGD optimizer, momentum of 0.9 and weight decay of 10e-4. We use a learning rate of 0.02 for all tasks. For MS COCO, we adopt AdamW as the optimizer, weight deacy of 0.01 and learning rate of 5e-5. Batch size is set to 16 for both datasets. In NSGP, we follow the adaptive selecting stategy proposed in  (Lu et al., 2024) to keep the singular vaules. We sample 9 extra fine-grained prototypes to complement the coarse prototype, 10 prototypes are used in total. The radius $r$ is set to 0.6. In our method, we incorporate a pseudo-labeling strategy to solve the foreground shift problem as implemented in BPF (Mo et al., 2024). All of our experiments were conducted on 2 RTX 3090 GPU.
@@ Appendix C Is the RoI Head robust to low-quality proposals?
Following  (Lu et al., 2024), we adaptively determine the nullity $R$ and retain $U_{t}^{\prime}$ correspond to $R$ smallest diagonal singular vaules $\lambda$ in $\Lambda_{t}$. Finally, the projection matrix for $(t+1)$-th training stage is obtained by
@@
It is a common practice in previous works (Wang et al., 2021; Lu et al., 2024) to normalize the $\cal B$ as

### v3

@@ Abstract
Catastrophic forgetting is a critical chanllenge for incremental object detection (IOD). Most existing methods treat the detector monolithically, relying on instance replay or knowledge distillation without analyzing component-specific forgetting. Through dissection of Faster R-CNN, we reveal a key insight: Catastrophic forgetting is predominantly localized to the RoI Head classifier, while regressors retain robustness across incremental stages. This finding challenges conventional assumptions, motivating us to develop a framework termed NSGP-RePRE. Regional Prototype Replay (RePRE) mitigates classifier forgetting via replay of two types of prototypes: coarse prototypes represent class-wise semantic centers of RoI features, while fine-grained prototypes model intra-class variations. Null Space Gradient Projection (NSGP) is further introduced to eliminate prototype-feature misalignment by updating the feature extractor in directions orthogonal to subspace of old inputs via gradient projection, aligning RePRE with incremental learning dynamics. Our simple yet effective design allows NSGP-RePRE to achieve state-of-the-art performance on the Pascal VOC and MS COCO datasets under various settings. Our work not only advances IOD methodology but also provide pivotal insights for catastrophic forgetting mitigation in IOD. Code is available at https://github.com/fanrena/NSGP-RePRE .
@@ 2.1 Incremental Learning for Classification
Most influential incremental learning studies have focused on classification tasks. Some regularization-based methods enforce the stability of logits (Li & Hoiem, 2017; Zhang et al., 2023; Yan et al., 2025) or intermediate features (Simon et al., 2021) to preserve the learned knowledge, while others apply restrictions on the weight of the model (Kirkpatrick et al., 2017) or on gradients during optimization (Lopez-Paz & Ranzato, 2017; Wang et al., 2021). Structure-based methods are dedicated to learning specific parameters for different tasks, with a dynamically expanding architecture (Rusu et al., 2016) or grouped parameters in a static model (Fernando et al., 2017). For replay-based methods, they can be divided into experience replay (Buzzega et al., 2020; Zhu et al., 2021; Kong et al., 2023) and generative replay methods (Zhai et al., 2020; Kemker & Kanan, 2017), depending on the examples stored in a buffer or generated with a model. Recently, incremental learning based on foundation models such as CLIP (Radford et al., 2021) has also attracted attention. Research works such as L2P (Wang et al., 2022), O-LoRA (Wang et al., 2023a), and VPT-NSP2 (Lu et al., ) attempt to learn continuously based on the parameter-efficient transfer learning technique (Zhou et al., 2022; Jia et al., 2022; Xing et al., 2023; Zhang et al., 2025) have achieved superior performance.
@@ 6 Conclusion
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant 62476223, 62176198, 62201467, the Key Research and Development Program of Shaanxi Province under Grant 2024GX-YBXM-135, in part by the Young Talent Fund of Xi’an Association for Science and Technology under Grant 959202313088, Innovation Capability Support Program of Shaanxi (No. 2024ZC-KJXX-043).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgements
Similar to previous works, we use the Faster R-CNN architecture with a Resnet-50 (He et al., 2016) backbone pre-trained in ImageNet (Russakovsky et al., 2015). On PASCAL VOC dataset, we train the network with SGD optimizer, momentum of 0.9 and weight decay of 10e-4. We use a learning rate of 0.02 for all tasks. For MS COCO, we adopt AdamW as the optimizer, weight deacy of 0.01 and learning rate of 5e-5. Batch size is set to 16 for both datasets. In NSGP, we follow the adaptive selecting stategy proposed in  (Lu et al., ) to keep the singular vaules. We sample 9 extra fine-grained prototypes to complement the coarse prototype, 10 prototypes are used in total. The radius $r$ is set to 0.6. In our method, we incorporate a pseudo-labeling strategy to solve the foreground shift problem as implemented in BPF (Mo et al., 2024). All of our experiments were conducted on 2 RTX 3090 GPU.
@@ Appendix C Is the RoI Head robust to low-quality proposals?
Following  (Lu et al., ), we adaptively determine the nullity $R$ and retain $U_{t}^{\prime}$ correspond to $R$ smallest diagonal singular vaules $\lambda$ in $\Lambda_{t}$. Finally, the projection matrix for $(t+1)$-th training stage is obtained by
@@
It is a common practice in previous works (Wang et al., 2021; Lu et al., ) to normalize the $\cal B$ as